### v1

# Symmetry breaking effects in pion couplings to constituent quarks currents

Abstract
Pseudoscalar and axial neutral and charged pion-constituent quark coupling constants are investigated with nondegenerate quark masses in different kinematical points, off shell and on shell pions and constituent quarks. For the neutral pion, mixing effects are introduced by means of the pion mixing to states $P_{0}$ and $P_{8}$, that give rise to the $\pi^{0}-\eta-\eta^{\prime}$ meson mixing, and mixing of quark currents via corresponding mixing interactions. The relative behavior of charged and neutral pion coupling constants to quarks may be nearly the same - in the framework of the constituent quark model - as the pion-nucleon coupling constants if mixings are introduced. A very small pion coupling to strange quark current is also obtained. The dependence of the positive and negative pion-constituent quark coupling constant on the non-degeneracy of quark masses, for emission and absorption processes, is identified.
1 Introduction
Pions, as the (quasi)Goldsone bosons of Strong Interactions, have small masses as compared to baryons. As a consequence, pion couplings to other hadrons, especially to nucleons, have importance for a large number of low and intermediary energy processes. In spite of their small masses, they are well described in the framework of the constituent quark model (CQM) [1] in which a large binding energy is required. In part of these models, dynamical chiral symmetry breaking (DChSB) plays a role of one of the mechanisms of mass generation and contributes for large quark (effective) masses, although other mechanisms may contribute [2]. In most versions of the CQM, mesons interact with constituent quarks belonging to a baryon. These couplings must lead to the measurable meson-baryon coupling constants that parameterize such interactions phenomenologically. Several of the corresponding coupling constants might follow well known relations usually based in chiral symmetry and its dynamical breakdown such as Goldberger Treiman (GT) and partially conserved axial current relations [3, 4]. With the improvement of theoretical models and approaches and the increase of experimental precision, it became important to understand further effects of isospin and flavor symmetry breakings (ISB and FSB).
There are several possible pion couplings to nucleons (or constituent quarks) with different Lorentz structures among which the axial and the pseudoscalar couplings are usually the most relevant ones. The corresponding coupling constants, and form factors, parameterize processes such as pion emission or absorption, besides being responsible for the Yukawa meson exchange among other processes. Isospin symmetry breaking effects in the pseudoscalar pion coupling to the nucleon have been investigated recently within different approaches [5, 6] leading to very similar conclusions that these effects must be very small. This is consistent with the fact that isospin symmetry breaking in flavor SU(2) QCD is ruled by up and down quark mass difference that is very small [7, 8, 9]. With different definitions of coupling constants, usually denoted by $f_{\pi^{a}NN^{\prime}}$ for $a=0,\pm$ and $N,N^{\prime}=n,p$, similar discrepancies between values of charged and neutral pion coupling constants were found, such as [6]: $g_{\pi^{0}nn}\simeq 13.397$, $g_{\pi^{0}pp}\simeq 13.227$ and $g_{\pi^{\pm}np}\simeq 13.226$. The pseudoscalar charged pion coupling constant is slightly smaller than the neutral one. In this work, the axial and pseudoscalar pion-constituent quark coupling constants will be derived from a quark determinant and investigated by considering non-degenerate up and down quark masses. Although the problem of establishing relations between quark and hadron level quantities might be firstly envisaged by means of group theoretical framework, the whole picture becomes more intricate due to confinement and to the dynamical character of Bethe-Salpeter and Faddeev type equations for bound quarks. Although this whole project may depend on the confinement mechanism and related properties, some aspects of hadron are seemingly independent of it according to phenomenology in which the CQM finds support for many hadrons. After all, color is confined, but flavor is not. With this program, the validity of the CQM description should be better defined and refined by investigating its consequences such as fine tuning (ISB or FSB) effects in meson couplings.
In several versions of the CQM, the pion axial coupling is usually considered to be $g_{A}=1$, including in large Nc EFT proposed by Weinberg [10, 11], being the axial charge directly related to the nucleon spin [12]. Recently, a quark determinant was analyzed in the presence of local meson fields and constituent (background) quark currents [11, 13, 14]. In a large quark mass expansion, both pseudoscalar and axial couplings appear, besides other leading scalar and vector couplings with the possibility of higher order (non-leading) effective couplings. In this approach, by starting with a leading term of the QCD effective quark action for quark-antiquark interaction, by means of a (non-perturbative) gluon exchange, quark currents are split into quantum quark currents and background quark currents, that are dressed by components of a non perturbative gluon propagator leading to constituent quark currents. Quantum quark currents lead to meson fields, and to the chiral condensate, by means of the auxiliary field method. Mesons structure and dynamics arise in flavor multiplets of quark-antiquark components similarly to previous developments [15, 16, 17, 18]. In this approach, pseudoscalar meson mixings due to FSB have been identified [19] being possible to add to other known mechanism for meson mixing by means of instanton-induced interactions [20]. In the present work, flavor mixing is investigated in two ways, by means of mixings for $\pi^{0}-\eta-\eta^{\prime}$ mesons and flavor mixing effective quark-antiquark interactions, $G_{ij}(\bar{\psi}\Gamma\lambda_{i}\psi)(\bar{\psi}\Gamma\lambda_{j}\psi)$ for $\Gamma=i\gamma_{5}$, derived recently by a similar method [21, 22].
Therefore, in the present work, an investigation of ISB (and FSB) in constituent quark-pion interactions is proposed by means of the pseudoscalar and axial pion couplings to constituent quark currents previously analyzed in the degenerate masses limit [11, 13]. Although the relation of these quark level couplings to the corresponding nucleon (or baryon in general) is not fully unambiguous for the reasons made explicit above, these quark level calculations and results should lead to more fine and precise assessment of to what extent, why and how constituent quark model works. Along these lines, the absolute values of the resulting coupling constants are not exactly the most relevant or interesting point of this work, but rather the effects of the nondegenerate quark masses in each channel, i.e. neutral or charged pions couplings to up or down quark currents, either pseudoscalar or axial. This might contribute to settle a route for analyzing hadron observable world with more complete grounds in low energy QCD. This work is organized as follows. In the next section some schematic explanation of the quark determinant in the presence of local meson fields and background quark currents, surrounded by a gluon cloud, is presented. The pseudoscalar and axial coupling constants are derived with the corresponding match equations. In Section (3) the possible mixings are analyzed for both pseudoscalar meson states and quark currents. For this, some mixing interactions obtained from flavor symmetry breaking in quark polarization [21, 22] are also considered. In Section (4) numerical results for the charged and neutral pion pseudoscalar and axial coupling constants are presented for different kinematical points: on shell and off shell (or massless) pions and constituent quarks. Some strangeness degree of freedom is shown to provide a small coupling to the neutral pion due to the mixings. In the last Section there is a Summary with conclusions.
2 Pion couplings from the quark determinant
The general procedure for the derivation of the quark determinant in the presence of background quark currents, surrounded by a gluon cloud represented by components of a gluon propagator, and local auxiliary meson fields displayed in quark-antiquark flavor U(3) nonets was developed in Refs. [11, 13, 14]. By selecting the (chiral) scalar and pseudoscalar auxiliary fields and by performing an usual chiral rotation, the scalar field nonet is eliminated in favor of non-linear parametrization of the (quasi) Goldstone boson octet as fluctuations around the vacuum. It can be written in terms of: $U=e^{\frac{iP\cdot\lambda}{F}}$ and $U^{\dagger}=e^{\frac{-iP\cdot\lambda}{F}}$, where $\lambda_{i}$ are the flavor Gell-Mann matrices and the pseudoscalar meson fields $P_{i}$ have canonical normalization. The quark determinant can be written as:
$S_{eff}$ $=$ $-i\;Tr\;\ln\;\left\{i\left[(S_{0}^{f})^{-1}+F(P_{R}U+P_{L}U^{\dagger})+\sum_{\phi}a_{\phi}J_{\phi}\right]\right\},$ (1)
where $Tr$ stands for traces of all discrete internal indices and integration of space-time coordinates, $P_{R,L}=\frac{1}{2}(1\pm\gamma_{5})$ are the right-hand and left-hand chirality projectors, $S_{0}^{f}$ is the quark propagator with a constituent quark mass ($M_{f}$) that can be written as $S_{0}^{f}=(i\not{\partial}-M_{f}+i\epsilon)^{-1}$. The quark masses, $M_{f}$ for $f=u,d,s$, are (large) constituent quark mass, usually considered to be generated by DChSB, but that may involve other mechanisms of mass generation such as trace anomaly as identified in Ref. [23]. The leading background dressed (singlet color) quark currents are the following:
$\sum_{\phi}a_{\phi}J_{\phi}=K_{0}\left[R(x-y)(J_{S}^{i}(x,y)+i\gamma_{5}\lambda_{i}J_{PS}^{i}(x,y))-\frac{{R}^{\mu\nu}(x-y)}{2}\gamma_{\mu}\lambda_{i}(\gamma_{5}J_{\mu}^{V,i}(x,y)+\gamma_{5}J_{\mu}^{A,i}(x,y))\right],$ (2)
where $K_{0}=4\alpha_{g}^{2}/9$ for $\alpha_{g}$ encodes the running quark-gluon coupling constant that will be incorporated latter into the effective gluon propagator. The (non-local) quark currents were defined respectively as $J_{S}^{i}=J_{S}^{i}(x,y)=(\bar{\psi}\lambda^{i}\psi)$, $J_{PS}^{i}=J_{PS}^{i}(x,y)=(\bar{\psi}i\gamma_{5}\lambda^{i}\psi)$, $J_{V,i}^{\mu}=J_{V,i}^{\mu}(x,y)=(\bar{\psi}\gamma^{\mu}\lambda^{i}\psi)$ and $J_{A,i}^{\mu}=J_{A,i}^{\mu}(x,y)=(\bar{\psi}\gamma^{\mu}\gamma_{5}\lambda^{i}\psi)$. The background quark currents are dressed by components of the effective gluon propagator that, by assuming it is decomposed in transverse and longitudinal terms, can be written, in momentum space, as:
$R(k)$ $=$ $3R_{T}(k)+R_{L}(k),$
$R^{\mu\nu}(k)$ $=$ $g^{\mu\nu}(R_{T}(k)+R_{L}(k))+2\frac{k^{\mu}k^{\nu}}{k^{2}}(R_{T}(k)-R_{L}(k)).$ (3)
Although all the pseudoscalar nonet have been included above, only the components $i=1,2,3$ for the pion field will be worked out, with some further attention to the components $i=0,8$ that undergo explicit flavor meson mixing with the neutral pion ($\pi^{0}-\eta-\eta^{\prime}$) and, correspondingly, flavor mixing of the quark currents.
This determinant can be treated in different ways. Along the same lines of previous works, the following dimensionless quantites will be considered in the determinant:
$X=S_{0}F(P_{R}U+P_{L}U^{\dagger}),\;\;\;\;\;\;Y=S_{0}R(x-y)J_{\phi}(x.y).$
These quantities suggest a large quark (and gluon) effective mass(es) expansion whose validity is not directly proved due to the dynamical character of the involved quantities. However, it is interesting to note that, this same expansion for the NJL-model (instead of quark-antiquark interaction mediated by gluon exchange) provide corrections of about $20\%-30\%$ for the NJL coupling constant [24, 25] - values within the perturbative regime. Also, it is interesting to note that, to provide a qualitative assessment of this expansion, the scalar quark current can replaced by an averaged value in the vacuum (rather taken its modulus as representative) that are the quark-antiquark scalar chiral condensate, $|<\bar{q}q>|<|(-250MeV)^{3}|$. Large (quark and gluon) effective masses for relatively low kinetic energies (phenomenologically, it can be considered to be below QCD cutoff) yield, for the propagators of internal lines, $S_{f}\sim 1/M_{f}$ and $R\sim 1/M_{g}^{2}$ (where $M_{g}$ is an effective gluon mass that might parameterize the gluon propagator). With such simple ad hoc consideration, by taking typical phenomenological values for these masses, $M_{f}\sim 350$MeV and $M_{g}\sim 400$ MeV, a typical term of this expansion can be written as
$Y\sim Y_{vac}\sim\frac{|<\bar{q}q>|}{M_{f}M_{g}^{2}}\sim 0.28,$
that is not largely modified with addition numerical factors and it is inside the corresponding perturbative regime. Moreover, it is worth noting that this expansion leads to all interaction terms allowed by the involved symmetries, along the lines of an effective field theory (EFT). As a matter of fact, the resulting leading pion couplings are the ones of the Weinberg’s large Nc EFT [10] with an additional scalar coupling (explicit symmetry breaking coupling) as shown in [11, 13]. The expansion of the whole determinant lead to three different sectors. Firstly, the meson sector that has bee exploited in many works in past decades for different versions of local and non-local quark interactions, and, specifically for the pion field, it yields Chiral Perturbation theory, for few works see [17, 16, 15, 18]. The quark currents sector lead to effective constituent quark interactions whose local limit has been exploited recently in some works and provide corrections to models such as the NJL model. Finally, the meson-constituent quark interactions sector that is exploited in the local limit for pions in the present work.
Eq. (1) can be worked out, except for an irrelevant constant, as
$S_{eff}^{B}=-i\;Tr\;\ln\;\left\{[1+X+Y]\right\}=-\frac{i}{2}Tr\;\ln\;\left\{[1+X+Y][1+X^{*}+Y^{*}]\right\}.$ (4)
The diagrammatic interpretation of the corresponding leading terms of the meson-quark interaction sector are shown in Fig. (1) and these diagrams - with one pion external leg - hold for both the pseudoscalar and the axial couplings with different form factors/coupling constants. Note that the flavor structure of the positive pion absorption coupling is different from the negative pion absorption but it is the same as the negative pion emission. In the expansion of the determinant, there also appear the scalar and vector (two-)pion couplings [13] that are of the same order of magnitude of the pseudoscalar and axial ones, although they present a considerably more intricate isospin structure and they will not be considered in this work.
Figure 1: Diagrammatic interpretation of the leading (charged and neutral) pion couplings to constituent quark currents (absorption) at one loop level. Pion and quark momenta are, respectively, $Q$ and $K$. In Diagram (c) $q=u,d$. Wiggly lines with a dot are components of a (non-perturbative) gluon (effective) propagator.
Although resulting interaction are non-local, by assuming large quark and gluon effective masses, it is possible to resolve the corresponding couplings in the local limit. Accordingly, they can be written as:
$I_{A}$ $=$ $G_{ij}^{ps}P_{i}J^{ps}_{j}+G_{ij}^{A}(\partial_{\mu}P_{i})J^{\mu}_{A,j},$ (5)
being that the (dimensionless) coupling constants are, in fact, form factors shown below. Alternatively, by considering Eq. (4), the expansion of the determinant leads to different channels of the couplings above. The corresponding terms can be written (with B) as:
$I^{B}$ $=$ $G_{ij}^{ps}P_{i}^{*}J^{ps}_{j}+G_{ij}^{A}(\partial_{\mu}P_{i}^{*})J^{\mu}_{A,j}.$ (6)
These coupling constants are in fact form factors that are defined according to the processes of Fig. (1). By assuming an effective gluon propagator, that will be given in Eq. (20), these form factors, for constant quark effective masses, can be defined as:
$G^{ps}_{ij}(K^{2},Q^{2},K\cdot Q)$ $=$ $N_{c}\;Tr_{F,D}\int\frac{d^{4}k}{(2\pi)^{4}}S_{0}^{f_{1}}(k+K)i\gamma_{5}\lambda_{i}S_{0}^{f_{2}}(k+K+Q)\lambda_{j}i\gamma_{5}R(-k),$ (7)
$G_{ij}^{A}(K^{2},Q^{2},K\cdot Q)g^{\mu\nu}$ $=$ $N_{c}F\;M_{f_{2}}\;Tr_{F,D}\int\frac{d^{4}k}{(2\pi)^{4}}\tilde{S}_{0}^{f_{1}}(k+K)\gamma^{\mu}\gamma_{5}\lambda_{i}\tilde{S}_{0}^{f_{2}}(k+K+Q)\lambda_{j}\gamma^{\nu}\gamma_{5}R(-k),$
where $Tr_{F,D}$ is the trace in flavor and Dirac indices, $\tilde{S}_{0}^{f}(k)=(k^{2}-M_{f}^{2}+i\epsilon)^{-1}$. Although only the states $i,j=1,2,3$ will be worked out, some other states are related by mixings, $i,j=0,8$. For the definitions above, in the case of nondegenerate quark mass $G_{f_{1}\neq f_{2}}=G_{f_{2}\neq f_{1}}$. The difference is, however, very small being proportional to the up and down quark mass differences in the isospin subgroup. Due to U(1) electromagnetic invariance and CP, one has $G_{ij}=G_{ji}$ and $G_{11}=G_{22}$. Some complementary definitions that will be useful to work out mixings are presented in the Appendix (A). The positive and negative pion form factors present exactly the same form factors since $G_{11}$ is given by Eq. (A.1). In fact, it involves the average of positive or negative pion absorption and emission processes.
3 Quark currents, meson states and mixings
By writing interactions (6) explicitly in terms of charge eigenstates, for the pseudoscalar coupling and corresponding pion form factors, it yields:
$I_{ps}^{B}$ $=$ $G_{11}^{ps}(\pi^{+}\bar{d}i\gamma_{5}u+\pi^{-}\bar{u}i\gamma_{5}d)+G_{33}^{ps}\pi_{3}(\bar{u}i\gamma_{5}u-\bar{d}i\gamma_{5}d),$ (8)
And similarly for the axial couplings. The charged pion couplings can be written as:
$I_{\pi^{\pm}}^{ps}=(G_{ud}^{ps}+G_{du}^{ps})(\pi^{+}\;\bar{d}i\gamma_{5}u+\pi^{-}\;\bar{u}i\gamma_{5}d),$ (9)
where $G_{f_{1}f_{2}}$ are defined in Eqs. (A.1)-(A.7). Note that these couplings correspond to an average of emitted and absorbed positive or negative pion. The separate coupling for emission or absorbed pion are equal in the limit of degenerate quark masses.
3.1 Mixings
The complete pseudoscalar interaction terms can be written, for the pion, i.e. $i,j=1,2,3$ and the corresponding mixing terms, $(i,j)=(0,3)$ or $(i,j)=(3,8)$, as:
$I_{ps}^{mix-ij}\equiv G_{ij}^{ps}P_{i}J_{j}^{ps}=G_{11}^{ps}P_{1}J_{1}^{ps}+G_{22}^{ps}P_{2}J_{2}^{ps}+G_{33}^{ps}P_{3}J_{3}^{ps}+G_{38}^{ps}P_{3}J_{8}^{ps}+G_{30}^{ps}P_{3}J_{0}^{ps},$ (10)
Being the flavor structure of the axial couplings analogous. For the neutral pion, if the mixing terms in Eq. (10) are not taken into account, the neutral pion coupling becomes the following one:
$I_{\pi^{0}}^{nomixing}=(G_{uu}^{ps}+G_{dd}^{ps})P_{3}\;\left(\bar{u}i\gamma_{5}u-\bar{d}i\gamma_{5}d\right)=G_{ps}^{\pi^{0}}(0)\;P_{3}\;\bar{\psi}\lambda_{3}i\gamma_{5}\psi.$ (11)
From this equation the following coupling constant, without mixing, can be defined $G_{ps}^{\pi^{0}}(0)$), and similarly to the axial coupling constant. If the explicit mixing terms between the neutral (either meson or current) states $i,j=0,3,8$ are considered, by writing explicitly the interactions, the charged and neutral pion couplings can be rewritten as:
$I_{\pi^{0}}^{mixing}$ $=$ $G_{uu}^{ps}P_{3}\;\bar{u}i\gamma_{5}u-G_{dd}^{ps}P_{3}\;\bar{d}i\gamma_{5}d.$ (12)
Therefore the neutral pion couples differently to up or down quark currents. However, the neutral pion coupling to a constituent quark current might be, still, interpreted as an average of the up and down couplings so that Eq. (11) seems reasonable in the presence of these mixings, i..e
$I_{\pi^{0}}^{mixing}\sim G^{ps}_{\pi^{0}}(ij)P_{3}\;(\bar{\psi}i\gamma_{5}\lambda_{3}\psi).$ (13)
Another mixing term can be taken into account. For instance, there might appear neutral pions from to the oscillations of $\eta,\eta^{\prime}$ due to meson mixings. There are two different possible ways meson mixing can show up. First in processes that lead to pions, etas and etaprimes for which contamination from $\eta,\eta^{\prime}$ oscillations can lead to contributions for the neutral pion coupling to constituent quarks/nucleon. Second, in processes in which the original neutral pion oscillates to $\eta$ or $\eta^{\prime}$ that can couple to quarks/nucleons. Consider the following parametrization of these meson mixings (rotations):
$P_{0}$ $=$ $\bar{G}_{80}\eta+\bar{G}_{03}\pi^{0}+a_{0}\eta^{\prime}$
$P_{8}$ $=$ $\bar{G}_{08}\eta^{\prime}+\bar{G}_{83}\pi^{0}+a_{8}\eta,$
$P_{3}$ $=$ $a_{3}\pi^{0}+\bar{G}_{38}\eta+\bar{G}_{30}\eta^{\prime},$ (14)
where the pion receives the smallest mixings and then $a_{3}<1$ with $a_{3}\sim 1$. These equations have been inspired in usual parametrization [26, 27]. Only two mixing terms will be needed below, $\eta-\pi^{0}$ or $\eta^{\prime}-\pi^{0}$ by means of $\bar{G}_{03}$ and $\bar{G}_{38}$. Furthermore, two types of interactions will be analyzed, the pseudoscalar and axial, so that all the coupling constants and mixing parameters will be labeled with an extra index $\xi=A,ps$ below. As a consequence, the following neutral pion coupling to quark currents, $J_{u}^{\xi},J_{d}^{\xi},J_{s}^{\xi}$, in both axial and pseudoscalar channels, are obtained:
$L_{\pi^{0}-J}^{\xi}$ $=$ $(\partial_{\xi}\pi^{0})a_{3}\left[J_{u}^{\xi}\left(G_{uu}^{\xi}+\frac{\bar{G}_{38}^{\xi}G_{uu}^{\xi}}{a_{3}}+\frac{2\bar{G}_{03}^{\xi}G_{uu}^{\xi}}{\sqrt{3}a_{3}}\right)-J_{d}^{\xi}\left(G_{dd}^{\xi}-\frac{\bar{G}_{38}^{\xi}G_{dd}^{\xi}}{a_{3}}-\frac{2\bar{G}_{03}^{\xi}G_{dd}^{\xi}}{\sqrt{3}a_{3}}\right)\right.$ (15)
$+$ $\left.\sqrt{2}J_{s}^{\xi}G_{ss}^{\xi}\left(\bar{G}_{38}^{\xi}-\frac{\bar{G}_{03}^{\xi}}{\sqrt{3}}\right)\right]$
$=$ $(\partial_{\xi}\pi^{0})a_{3}\left[G_{mix,uu}^{\xi}J_{u}^{\xi}-G_{mix,dd}^{\xi}J_{d}^{\xi}+G_{mix,ss}^{\xi}J_{s}^{\xi}\right],$
where:
$\mbox{for}\;\;\;\;\xi=ps$ ::: $(\partial_{\xi}\pi^{0})\to\pi^{0},\;\;\;\;\;\;\;\;J^{\xi}=J^{ps},$
$\mbox{for}\;\;\;\;\;\xi=A$ ::: $(\partial_{\xi}\pi^{0})\to\partial_{\mu}\pi^{0},\;\;\;\;\;J^{\xi}=J^{\mu}_{A}.$ (16)
Note the last term is a neutral pion coupling to the strange quark (axial or pseudoscalar) current. In the last line of Eq. (15), the neutral pion coupling to the different quark currents incorporated the mixings that lead to the neutral pion coupling to the strange quark current, $G^{\xi}_{mix,ss}$. The neutral pion couplings are corrected differently for $\bar{u}u$ and $\bar{d}d$ states, i.e. the corresponding corrections have opposite signs. By resorting to an analogy with the decomposition for Eq. (11) that yielded Eq. (12), the following neutral pion coupling to constituent quarks will be defined by using the definitions (3.1):
$I_{\pi^{0},mix}^{mix}$ $=$ $(\partial_{\xi}\pi^{0})a_{3}(G_{mix,uu}^{\xi}(ij)+G_{mix,dd}^{\xi}(ij))J^{\xi}_{3}.$ (17)
The meson mixing parameters (3.1) will be extracted from a simple association of meson mixing interactions $G_{i\neq j}$ that lead to diagonal meson states by starting from flavor eigenstates by considering separate diagonalization/procedure for each of the mixing interactions [21, 22]. This, nearly ad hoc, prescription yields:
$g_{03}^{\xi}\equiv\bar{G}_{03}^{\xi}=\frac{G_{03}^{\xi}}{2(G_{00}^{\xi}+G_{33}^{\xi})},\;\;\;\;\;g_{38}^{\xi}\equiv\bar{G}_{38}^{\xi}=\frac{G_{38}^{\xi}}{2(G_{33}^{\xi}+G_{88}^{\xi})}.$ (18)
If these mixings were simply $\bar{G}_{03}^{\xi}=G_{03}^{\xi}$ and $\bar{G}_{38}^{\xi}=G_{38}^{\xi}$ the pion coupling to the strange-quark current would disappear, i.e. $G_{mix,ss}^{\xi}=0$.
3.2 Renormalization condition
To provide numerical results, the following renormalization condition will be considered by fixing the pseudoscalar charged pion coupling constant with a value based in phenomenology [28] (off shell pion and on shell quark):
$G^{ps}_{\pi^{\pm}}(K^{2}=M^{2},Q^{2}=0)=13.$ (19)
It turns out that, for the sets of parameters considered in this work, results will provide nearly $g_{A}(M^{2},0)^{\pi^{\pm}}\sim 1$ without further conditions. A different condition could be imposed by assuming each quark of the nucleon to be responsible for part of the couplings to mesons, such that $G_{ps}\sim 13/3$. However, for the condition above, pion-nucleon coupling constants should arise from an average of pion-constituent quark coupling constants. This might be considered in the Goldberber-Treiman (GT)-relation if the nucleon mass is considered instead of the constituent quark mass ($M_{q}$). Otherwise, with a constituent quark mass, GT-relation should be written as $\frac{G^{ps}}{g^{A}}=\frac{M_{q}}{f_{\pi}}(1+\Delta_{GT})$ being $\Delta_{GT}$ not small with respect to 1. Conversely, in other works [13, 30], an alternative condition was considered, i.e. to fix either $g_{A}=1$ or $g_{v}=12$ (vector meson coupling to constituent quarks) for the CQM.
4 Numerical Results
The following longitudinal (confining) effective gluon propagator, inspired in [29] and tested in different works [30, 13], has been adopted:
$D^{\mu\nu}(k)$ $=$ $\frac{K_{F}g^{\mu\nu}}{(k^{2}-M_{g}^{2})^{2}},$ (20)
where the strong running coupling constant ($\alpha_{g}^{2}$) has been incorporated implicitely in $K_{F}$. It will be considered $M_{g}=500$MeV, and the normalization $K_{F}$ is determined by imposing the renormalization condition (19). It has been shown, in Refs. [30, 13], to produce similar numerical values and behavior to a different gluon propagator extracted from Schwinger Dyson equations [31] for spacelike meson-constituent quark form factors
Given that there is an intrinsic ambiguity in defining values for the quark effective masses, results will be shown for different choices of $M_{u},M_{d}$ (and also $M_{s}$). These masses, for the chosen gluon effective propagator, determine $G_{ij}^{\xi}$ (both $G_{ps}$ and $G_{A}$) numerically, and the different sets of parameters (s.o.p.) are shown in Table (1). Sets of parameters $A_{1}$ and $A_{2}$ were extracted from [21]. The third $A_{3}$ is based in often values found in the literature.
Table 1: Sets of parameters for quark effective masses.
set of $M_{u}$ $M_{d}$ $M_{s}$
parameters MeV MeV MeV
$A_{1}$ 389 399 600
$A_{2}$ 307 319 349
$A_{3}$ 340 350 550
$m_{\pi^{0}}$= 134.98 , $m_{\pi^{\pm}}$= 139.57 (MeV)
The mixing parameters defined in Eq. (3.1) were calculated at the same kinematical point ($K^{2}=M_{f}^{2},Q^{2}=0$) for each of the s.o.p. The relative difference from one kinematical point to another is much smaller. They are given by
$A_{1}$ $\bar{G}_{03}^{ps}=6.3\times 10^{-3},\;\;\;\;\;\bar{G}_{38}^{ps}=5.9\times 10^{-3},\;\;\;\;\;\bar{G}_{03}^{A}=-1.6\times 10^{-3},\;\;\;\;\;\bar{G}_{38}^{A}=-1.1\times 10^{-3},$
$A_{2}$ $\bar{G}_{03}^{ps}=3.9\times 10^{-3},\;\;\;\;\;\bar{G}_{38}^{ps}=2.8\times 10^{-3}\;\;\;\;\;\bar{G}_{03}^{A}=-2.7\times 10^{-3},\;\;\;\;\;\bar{G}_{38}^{A}=-1.9\times 10^{-3},$
$A_{3}$ $\bar{G}_{03}^{ps}=4.5\times 10^{-3},\;\;\;\;\;\bar{G}_{38}^{ps}=3.8\times 10^{-3}\;\;\;\;\;\bar{G}_{03}^{A}=-1.9\times 10^{-3},\;\;\;\;\;\bar{G}_{38}^{A}=-1.3\times 10^{-3},$ (21)
In Table (2) values of pseudoscalar coupling constants $G_{ps}$, defined in different kinematical points (off shell or on shell constituent quark and pion) are shown for the different sets of parameters presented in Table (1). Besides the charged pion coupling constant, the neutral pion coupling constant is exhibited with and without flavor states mixings, respectively $G^{ps}_{\pi^{0}}(0)$ and $G^{ps}_{\pi^{0}}(ij)$ from Eqs. (11) and (17). As shown above the neutral pion coupling can be decomposed in the couplings of states $\bar{u}u$ and $\bar{d}d$ states, that might also receive contributions from flavor state mixings - and these are shown in the middle part of this Table. In the last two columns the pseudoscalar coupling constant of the neutral pion coupling to strange quark current $G_{mix,ss}^{ps}$ is exhibited with a parameter $x_{ps}^{m}$ that will be defined below.
Results for s.o.p. $A_{2}$ ($A_{1}$), with the lower (larger) values of quark masses, present the smallest (largest) variation of values for the different kinematical points. The charged pion coupling constants are smaller than the neutral pion ones except for the case of the fourth kinematical point: off shell quarks and on shell pions $(K^{2}=0,Q^{2}=M_{\pi}^{2})$. The neutral pion coupling to the $\bar{u}u$ quark current has larger strength than the coupling to $\bar{d}d$ for all kinematical points, and for both cases, without and with mixings, $G^{ps}_{uu}(0)$ and $G^{ps}_{uu}(ij)$ respectively.
Recent analysis of the pion-nucleon pseudoscalar coupling constant calculated with different approaches [6, 5] exhibit a very similar trend with small numerical difference between charged and neutral pion couplings to nucleons. The neutral pion -nucleon coupling constants are slightly larger than the charged pion coupling coupling constant. For the sake of comparison of the behavior, the following values can be extracted from the calculation with the EFT approach [6]: $g_{\pi^{0}pp}=13.22688$, $g_{\pi^{0}nn}=13.39731$, $g_{\pi^{\pm}np}=13.22611$. According to the quark model, proton and neutron structures suggest that the corresponding coupling to the pion might be a sort of sum or average for the pion coupling to corresponding states, $p\sim uud$ and $n\sim ddu$ respectively. However, due to the choice of the renormalization condition ($G^{ps}_{\pi^{p}m}=13$), the resulting neutral pion coupling to nucleons, as a first simplified approach, might be considered as a simple average:
$g_{\pi^{0}nn}\sim\frac{2G_{\pi^{0}dd}+G_{\pi^{0}uu}}{3},\;\;\;\;\;\;\;\;g_{\pi^{0}pp}\sim\frac{2G_{\pi^{0}uu}+G_{\pi^{0}dd}}{3}.$ (22)
Concerning the charged pion coupling, the behavior shown in the Table is consistent with the results for the charged pion-neutron-proton and neutral pion-proton coupling constants such that that is $g_{\pi^{\pm}np}<g_{\pi^{0}pp},g_{\pi^{0}nn}$ and $g_{\pi^{\pm}np}\sim g_{\pi^{0}pp}$ [6, 5]. Along these lines, it is also expected that $g_{\pi^{0}nn}>g_{\pi^{0}pp}$. This hierarchy in pion-nucleon coupling constants, however, does not happen with the values shown in the Table for (22) for the model of (22), since it would require $G_{\pi^{0}dd}>G_{\pi^{0}uu}$ i.e. the opposite behavior to the one shown in the Table. It is shown below that mixings can mend this behavior.
By varying the strength of the mixing interactions by a constant uniform factor $x^{m}_{ps}$, the pseudoscalar coupling constants $G^{ps}_{uu}$ and $G^{ps}_{dd}$ can flip their hierarchy if $x_{ps}^{m}<0$. Therefore mixing parameters were redefined by:
$\bar{G}^{ps}_{i\neq j}\to x^{m}_{ps}\times\bar{G}^{ps}_{i\neq j}.$ (23)
This redefinition of the mixing parameters can include further mixing interactions, not only due to isospin/flavor symmetry breaking, i.e. they might include ’t Hooft interactions. The resulting values for $x_{ps}^{m}$ for which the coupling constants $G^{ps}_{uu}$ become smaller than $G^{ps}_{dd}$ are in the last column of the Table (2). However, with this ad hoc change in the mixing parameters to make quark interactions compatible with neutral pion coupling to nucleons the hierarchy of neutral and charged pion couplings may be inverted unless a redefinition of the charged pion coupling constant is done.
In Table (3) values of the axial $G_{A}$ coupling constant are shown for the same cases of Table (2). The behavior for $G_{A}$ at different kinematical points are different from the results of the $G_{ps}$, being the charged pion coupling constants slightly larger than the neutral pion ones. Although the variation of the axial coupling constant is smaller than the case of the pseudoscalar coupling constant of the previous Table, the relative deviations are nearly of the same order of magnitude. As discussed above, the strength of the pseudoscalar coupling constant is fixed with the renormalization condition (19) and this settles the values shown in the Table $G_{A}(K^{2},Q^{2})$. The s.o.p. $A_{2}$ (lower quark effective masses) presents the closer value of $G_{A}$ to the expected value for the CQM, $G_{A}=1$ [32, 10]. However, the author is not aware of previous investigations of the contribution of the isospin breaking for the pion axial coupling constant in the literature. In the last column of Fig. (3) the coupling constant for the axial pion - strange quark current is presented. The strength of this strange axial pion coupling constant is relatively smaller than the pseudoscalar one. The axial $G^{uu}_{A}$ are already weaker than $G^{dd}_{A}$ and the mixing interactions increase this difference.
Table 2: Pseudoscalar pion coupling constants for charged and neutral pions at different kinematical points, without ($G^{ps}_{\pi^{0}}(0)$) (11) and with mixing interactions $G_{i\neq j}$ ($G^{ps}_{\pi^{0}}(ij)$) (17). Values of pseudoscalar coupling constants for the $\pi^{0}$ with states $\bar{u}u$ and $\bar{d}d$ (shortened notation $G^{ps}_{uu}$ and $G^{ps}_{dd}$) also shown without $(0)$ and with $(ij)$ mixing interactions. In the two last columns, the pseudoscalar coupling constant of neutral pion -strange quark current interaction and the parameter used to vary mixing interactions $x_{ps}^{m}$ defined in Eq. (23).
s.o.p./$(K^{2},Q^{2})$ $G^{ps}_{\pi^{\pm}}$ $G^{ps}_{\pi^{0}}$ (0) $G^{ps}_{\pi^{0}}$ ($ij$) $G^{ps}_{uu}$ (0) $G^{ps}_{uu}$ ($ij$) $G^{ps}_{dd}$ (0) $G^{ps}_{dd}$ ($ij$) $G^{ps}_{mix,ss}$ $x^{m}_{ps}$
$A_{1}$
$(M_{f}^{2},0)$ 13.000 13.001 13.007 13.471 13.648 12.531 12.367 0.05 -1.7
$(M_{f}^{2},M_{\pi}^{2})$ 10.446 10.546 10.553 11.098 11.244 9.994 9.863 0.37 -1.2
$(0,0)$ 17.870 17.871 17.875 18.172 18.411 17.569 17.338 - 0.01 -2.3
$(0,M_{\pi}^{2})$ 18.106 18.091 18.095 18.399 18.642 17.782 17.545 -0.01 -2.3
$A_{2}$
$(M_{f}^{2},0)$ 13.000 13.000 13.003 13.363 13.463 12.637 12.543 -0.02 -2.6
$(M_{f}^{2},M_{\pi}^{2})$ 12.263 12.284 12.287 12.708 12.802 11.861 11.773 -0.02 -2.5
$(0,0)$ 14.550 15.550 14.552 14.818 14.928 14.283 14.177 -0.01 -2.6
$(0,M_{\pi}^{2})$ 14.808 14.792 14.794 15.069 15.181 14.514 14.406 -0.01 -2.2
$A_{3}$
$\;(M_{f}^{2},0)$ 13.000 13.000 13.004 13.352 13.473 12.649 12.535 0.004 -2.0
$(M_{f}^{2},M_{\pi}^{2})$ 11.735 11.779 11.783 12.191 12.301 11.368 11.266 0.04 -1.8
$(0,0)$ 15.458 15.459 15.461 15.704 15.846 15.213 15.076 -0.01 -2.3
$(0,M_{\pi}^{2})$ 15.702 15.686 15.689 15.940 16.083 15.433 15.294 -0.01 -2.3
Table 3: Axial pion coupling constants for charged and neutral pions at different kinematical points, without ($G^{A}_{\pi^{0}}(0)$) and with ($G^{A}_{\pi^{0}}(ij)$) mixing interactions $G_{i\neq j}$, similarly to Eqs. (11) and (17). Values presented for the the states $\bar{u}u$ and $\bar{d}d$ states without $(0)$ and with $(ij)$ mixing interactions. In the last column the axial coupling constant of neutral pion -strange quark current.
s.o.p./$(K^{2},Q^{2})$ $G^{A}_{\pi^{\pm}}$ $G^{A}_{\pi^{0}}$ (0) $G^{A}_{\pi^{0}}$ ($ij$) $G^{A}_{uu}$ (0) $G^{A}_{uu}$ ($ij$) $G^{A}_{dd}$ (0) $G^{A}_{dd}$ ($ij$) $G^{A}_{mix,ss}$
$A_{1}$
$(M_{f}^{2},0)$ 1.6027 1.6026 1.6026 1.6000 1.5954 1.6051 1.6099 0.0001
$(M_{f}^{2},M_{\pi}^{2})$ 1.8888 1.8750 1.8750 1.8717 1.8662 1.8784 1.8839 0.0002
$(0,0)$ 1.1936 1.1936 1.1935 1.1970 1.1935 1.1901 1.1936 - 0.0002
$(0,M_{\pi}^{2})$ 1.2230 1.2210 1.2210 1.2250 1.2214 1.2170 1.2206 - 0.0002
$A_{2}$
$(M_{f}^{2},0)$ 0.9736 0.9735 0.9736 0.9706 0.9656 0.9765 0.9815 0.0002
$(M_{f}^{2},M_{\pi}^{2})$ 1.1442 1.1357 1.1357 1.1320 1.1262 1.1394 1.1452 0.0002
$(0,0)$ 0.7801 0.7801 0.7801 0.7819 0.7779 0.7782 0.7822 -0.0001
$(0,M_{\pi}^{2})$ 0.8062 0.8045 0.8045 0.8070 0.8029 0.8019 0.8060 -0.0001
$A_{3}$
$(M_{f}^{2},0)$ 1.1602 1.1601 1.1601 1.1577 1.1536 1.1626 1.1666 0.0001
$(M_{f}^{2},M_{\pi}^{2})$ 1.3647 1.3547 1.3547 1.3517 1.3469 1.3578 1.3626 0.0002
$(0,0)$ 0.9030 0.9030 0.9029 0.9052 0.9020 0.9008 0.9040 -0.0001
$(0,M_{\pi}^{2})$ 0.9296 0.9278 0.9278 0.9305 0.9273 0.9251 0.9283 -0.0001
As noted before, the positive pion absorption (negative pion emission) coupling constant in both channels, $G_{du}^{\xi}(K^{2},Q^{2})$ ($\xi=ps,A$), might be different from the negative pion absorption (positive pion emission) coupling constant, $G_{ud}^{\xi}(K^{2},Q^{2})$ in the case of non-degenerate quark masses. The difference between these two charged pion coupling constants, for both pseudoscalar and axial couplings, is shown in Fig. (2), by means of the following (on shell quarks) ratios:
$D_{u-d}\equiv\frac{G_{ud}^{\xi}(K^{2}=M_{f}^{2},Q^{2})-G_{du}^{\xi}(K^{2}=M_{f}^{2},Q^{2})}{G_{ud}^{\xi}(K^{2}=M_{f}^{2},Q^{2})+G_{du}^{\xi}(K^{2}=M_{f}^{2},Q^{2})}.$ (24)
Two cases are exhibited, $Q^{2}=M_{\pi^{\pm}}^{2}$ and $Q^{2}=0$. It is seen that the variation is very small and basically linear, i.e. goes with the difference of masses $M_{d}-M_{u}$ in the leading order. The relative variation of the axial coupling constant is smaller than the pseudoscalar. The case of on shell or massive (off shell or massless) pion presents a larger variation for the pseudoscalar (axial) coupling constant.
Figure 2: Dependence of the difference of charged pion coupling constants (24) as a function of the mass ratio $M_{d}/M_{u}$, for the cases $Q^{2}=0$ and $Q^{2}=M_{\pi^{\pm}}^{2}$, with $K^{2}=M_{f}^{2}$.
4.1 Pion coupling to Strange currents
Both pion couplings to strangeness current, the pseudo-scalar and the axial ones, are very small and they depend on the strange quark effective mass. The behavior of the corresponding coupling constants will be shown with respect to the pseudoscalar and to the axial coupling constants (that are independent of $M_{s}$) by means of the following ratios with respect to the charged pion coupling constants with definitions in Eq. (15):
$G_{ss}^{ps}=\frac{G_{mix,ss}^{ps}}{\frac{G_{ud}^{ps}+G_{du}^{ps}}{2}},\;\;\;\;\;\;\;G_{ss}^{A}=\frac{G_{mix,ss}^{A}}{\frac{G_{ud}^{A}+G_{du}^{A}}{2}},$ (25)
where the coupling constants have been computed for on shell quarks $K^{2}=M_{f}^{2}$ and off shell pions $Q^{2}=0$. In Figs. (3) and (4) the ratios (25) are exhibited as functions of the strange quark effective mass that is freely varied for each of the s.o.p. The pseudoscalar pion coupling to the strangeness current is relatively larger than the axial pion coupling to the corresponding strange quark current. Besides that, $G_{mix,ss}^{ps}$ changes to negative values for $M_{s}>0.8$ GeV, for all the three s.o.p. For $G_{ss}^{ps}$ the s.o.p. with slightly different behavior is $A_{1}$ that has larger effective masses whereas the relative behavior of $G_{ss}^{A}$ is not really determined by the strength of the effective masses of each of the s.o.p.
Figure 3: Ratio $G_{ss}^{ps}$ defined in (25) as a function of $M_{s}$ for the three s.o.p.
Figure 4: Ratio $G_{ss}^{A}$ defined in (25) as a function of $M_{s}$ for the three s.o.p.
5 Summary and conclusions
Pseudoscalar and axial charged and neutral pion -constituent quark coupling constants were derived and investigated by considering quark mass non-degeneracy within an approach proposed previously by the author. These coupling constants were calculated for different sets of parameters of constituent quark masses and in different kinematical points: for on shell or off shell pions and constituent quarks. The pseudoscalar coupling constant calculated for off shell (or massless) charged pion and on shell constituent quark was fixed as a renormalization condition by identifying it to a phenomenological value $G^{ps}_{\pi^{\pm}}(K^{2}=M_{f}^{2},Q^{2}=0)=13$. With this condition, the resulting axial coupling constant is obtained with values close to the value adopted for the CQM, i.e. $g_{A}\sim 1$. Being all the coupling constants functions of the quark constituent masses (and not the small current quark masses), the Goldberger-Treiman relation at the quark level presents, in fact, a large correction for the values $G_{ps}=13$ and $g_{A}=1$. Results suggest that lower values of quark effective masses lead to smaller differences between the neutral and charged pion couplings (both axial and pseudoscalar). A very small difference between charged and neutral pion couplings has been noted in the pion-nucleon pseudoscalar coupling constant. The neutral pion coupling to quark currents $\bar{u}u$ and $\bar{d}d$ was compared to neutral pion coupling to protons and neutrons (respectively) and this comparison suggests mixing interactions stronger than those provided by flavor symmetry breaking are needed. This could be provided by instanton induced ’t Hooft-like interactions.
The resulting charged pion coupling constants from the quark determinant present full U(1) and CP invariance with $G_{11}=G_{22}$. It corresponds to an average of emitted and absorbed (positive or negative) pion, being therefore the same for each charge eigenstate or for each type of process (emission or absorption). For non-degenerate quark masses, there appears a small difference between absorption and emission of a positive or negative pion. This difference is proportional - in the leading order - to the quark mass difference $(M_{d}-M_{u})$. However, high precision experimental results/fittings would be needed to identify such difference.
Interactions that mix neutral flavor eigenstates $i,j=0,3,8$, leading to usual $\pi^{0}-\eta-\eta^{\prime}$ mixing, have been taken into account for both the quark currents and for the neutral pion. The shifts in the neutral pion coupling to the quark current $\bar{u}u$ are opposite to the shifts in the coupling to the current $\bar{d}d$. Besides that, a quite weak neutral pion coupling to the strange current $\bar{s}s$ was obtained for both axial and pseudoscalar cases. The relative strength of these couplings to strangeness, with respect to the pseudoscalar and axial ones, is very small $\sim 10^{-3},10^{-4}$, so that their contributions for neutral pion couplings to strange baryons would receive very tiny shifts only possibly identified with high precision measurements. Although the hierarchy of the strength of the flavor dependent coupling constants might nearly follow the one from the nucleon-pion couplings, the behavior of these coupling constants by inserting these couplings into the three-quark bound system for the nucleon can bring further insights into this subject. Besides that, at finite temperature, baryon density or external electromagnetic fields, all these couplings must receive non-negligible contributions. This type of investigation might seem not appropriated for the understanding of hadron interactions because quarks are confined in hadrons. However it is important to emphasize that confinement operates in color degrees of freedom whereas flavor (isospin) cannot be (completely) confined, otherwise all the quark-based description of hadrons, and maybe QCD, would be invalidated.
Acknowledgements
The author thanks a F.L.B. is a member of INCT-FNA, Proc. 464898/2014-5 and he acknowledges partial support from CNPq-312750/2021 and CNPq-407162/2023-2.
Appendix A Some relations
The following flavor-structures compose the coupling constants/form factors $G^{\xi}_{ij}$ in (7)
$G_{11}^{\xi}$ $=$ $(G_{ud}^{\xi}+G_{du}^{\xi}),$ (A.1)
$G_{33}^{\xi}$ $=$ $(G_{uu}^{\xi}+G_{dd}^{\xi}),$ (A.2)
$G_{03}^{\xi}$ $=$ $\sqrt{\frac{2}{3}}(G_{uu}^{\xi}-G_{dd}^{\xi}),$ (A.3)
$G_{08}^{\xi}$ $=$ $\sqrt{\frac{2}{3}}(G_{uu}^{\xi}+G_{dd}^{\xi}-2G_{ss}^{\xi}),$ (A.4)
$G_{38}^{\xi}$ $=$ $\frac{1}{\sqrt{3}}(G_{uu}^{\xi}-G_{dd}^{\xi}),$ (A.5)
$G_{00}^{\xi}$ $=$ $\frac{2}{3}(G_{uu}^{\xi}+G_{dd}^{\xi}+G_{ss}^{\xi}),$ (A.6)
$G_{88}^{\xi}$ $=$ $\frac{1}{3}(G_{uu}^{\xi}+G_{dd}^{\xi}+4G_{ss}^{\xi}).$ (A.7)

### v2

@@
-Symmetry breaking effects in pion couplings to constituent quarks currents
+Symmetry breaking effects in pion couplings to constituent quark currents
 Abstract
 Pseudoscalar and axial neutral and charged pion-constituent quark coupling constants are investigated with nondegenerate quark masses in different kinematical points, off shell and on shell pions and constituent quarks.
+By considering a large quark mass expansion of a quark determinant in the presence of local pion field and of constituent quark background currents, gluonic effects are considered by means of an effective gluon propagator that dresses quark currents.
 For the neutral pion, mixing effects are introduced by means of the pion mixing to states $P_{0}$ and $P_{8}$, that give rise to the $\pi^{0}-\eta-\eta^{\prime}$ meson mixing, and mixing of quark currents via corresponding mixing interactions.
@@
 As a consequence, pion couplings to other hadrons, especially to nucleons, have importance for a large number of low and intermediary energy processes.
-In spite of their small masses, they are well described in the framework of the constituent quark model (CQM) [1] in which a large binding energy is required.
+In spite of their small masses, they are well described in the framework of the constituent quark model (CQM) [1, 2] in which a large binding energy is required.
 In part of these models, dynamical chiral symmetry breaking (DChSB) plays a role of one of the mechanisms of mass generation and contributes for large quark (effective) masses, although other mechanisms may contribute [3].
@@
 These couplings must lead to the measurable meson-baryon coupling constants that parameterize such interactions phenomenologically.
-Several of the corresponding coupling constants might follow well known relations usually based in chiral symmetry and its dynamical breakdown such as Goldberger Treiman (GT) and partially conserved axial current relations [3, 4].
+Several of the corresponding coupling constants might follow well known relations usually based in chiral symmetry and its dynamical breakdown such as Goldberger Treiman relation (GTR) and partially conserved axial current [4, 5].
 With the improvement of theoretical models and approaches and the increase of experimental precision, it became important to understand further effects of isospin and flavor symmetry breakings (ISB and FSB).
-There are several possible pion couplings to nucleons (or constituent quarks) with different Lorentz structures among which the axial and the pseudoscalar couplings are usually the most relevant ones.
+There are several possible pion couplings to nucleons (or constituent quarks) with differerent Lorentz structures among which the axial and the pseudoscalar couplings are usually the most relevant ones.
 The corresponding coupling constants, and form factors, parameterize processes such as pion emission or absorption, besides being responsible for the Yukawa meson exchange among other processes.
-Isospin symmetry breaking effects in the pseudoscalar pion coupling to the nucleon have been investigated recently within different approaches [5, 6] leading to very similar conclusions that these effects must be very small.
-This is consistent with the fact that isospin symmetry breaking in flavor SU(2) QCD is ruled by up and down quark mass difference that is very small [7, 8, 9].
-With different definitions of coupling constants, usually denoted by $f_{\pi^{a}NN^{\prime}}$ for $a=0,\pm$ and $N,N^{\prime}=n,p$, similar discrepancies between values of charged and neutral pion coupling constants were found, such as [6]: $g_{\pi^{0}nn}\simeq 13.397$, $g_{\pi^{0}pp}\simeq 13.227$ and $g_{\pi^{\pm}np}\simeq 13.226$.
-The pseudoscalar charged pion coupling constant is slightly smaller than the neutral one.
-In this work, the axial and pseudoscalar pion-constituent quark coupling constants will be derived from a quark determinant and investigated by considering non-degenerate up and down quark masses.
-Although the problem of establishing relations between quark and hadron level quantities might be firstly envisaged by means of group theoretical framework, the whole picture becomes more intricate due to confinement and to the dynamical character of Bethe-Salpeter and Faddeev type equations for bound quarks.
-Although this whole project may depend on the confinement mechanism and related properties, some aspects of hadron are seemingly independent of it according to phenomenology in which the CQM finds support for many hadrons.
-After all, color is confined, but flavor is not.
-With this program, the validity of the CQM description should be better defined and refined by investigating its consequences such as fine tuning (ISB or FSB) effects in meson couplings.
+Isospin symmetry breaking effects in the pseudoscalar pion coupling to the nucleon have been investigated recently within different approaches [6, 7] leading to similar conclusions that these ISB effects must be very small.
+According to these works, the pseudoscalar neutral pion-neutron coupling constant is larger than the pseudoscalar neutral pion-proton coupling constant.
+However, it may be either slightly larger or considerably smaller than the charged pion coupling constant.
+Nevertheless, the resulting overall small ISB is consistent with the fact that isospin symmetry breaking in flavor SU(2) QCD is ruled by up and down quark mass difference that is very small [8, 9, 10].
+Other related consequences have been exploited in the literature, for example in [6, 11].
 In several versions of the CQM, the pion axial coupling is usually considered to be $g_{A}=1$, including in large Nc EFT proposed by Weinberg [12, 13], being the axial charge directly related to the nucleon spin [14].
+In the CQM, meson couplings to baryons, for example, are decomposed in meson couplings to constituent quarks [15].
 Recently, a quark determinant was analyzed in the presence of local meson fields and constituent (background) quark currents [13, 16, 17].
 In a large quark mass expansion, both pseudoscalar and axial couplings appear, besides other leading scalar and vector couplings with the possibility of higher order (non-leading) effective couplings.
-In this approach, by starting with a leading term of the QCD effective quark action for quark-antiquark interaction, by means of a (non-perturbative) gluon exchange, quark currents are split into quantum quark currents and background quark currents, that are dressed by components of a non perturbative gluon propagator leading to constituent quark currents.
+In this approach, by starting with a leading term of the QCD effective quark action for quark-antiquark interaction, by means of a (nonperturbative) gluon exchange, quark currents are split into quantum quark currents and background quark currents, that are dressed by components of a non perturbative gluon propagator leading to constituent quark currents.
 Quantum quark currents lead to meson fields, and to the chiral condensate, by means of the auxiliary field method.
@@
 In this approach, pseudoscalar meson mixings due to FSB have been identified [22] being possible to add to other known mechanism for meson mixing by means of instanton-induced interactions [23].
-In the present work, flavor mixing is investigated in two ways, by means of mixings for $\pi^{0}-\eta-\eta^{\prime}$ mesons and flavor mixing effective quark-antiquark interactions, $G_{ij}(\bar{\psi}\Gamma\lambda_{i}\psi)(\bar{\psi}\Gamma\lambda_{j}\psi)$ for $\Gamma=i\gamma_{5}$, derived recently by a similar method [21, 22].
-Therefore, in the present work, an investigation of ISB (and FSB) in constituent quark-pion interactions is proposed by means of the pseudoscalar and axial pion couplings to constituent quark currents previously analyzed in the degenerate masses limit [11, 13].
-Although the relation of these quark level couplings to the corresponding nucleon (or baryon in general) is not fully unambiguous for the reasons made explicit above, these quark level calculations and results should lead to more fine and precise assessment of to what extent, why and how constituent quark model works.
+In the present work, flavor mixing are investigated in two ways, by means of mixings for $\pi^{0}-\eta-\eta^{\prime}$ mesons and flavor mixing at the level of quark flavor currents in effective quark-antiquark interactions, $G_{ij}(\bar{\psi}\Gamma\lambda_{i}\psi)(\bar{\psi}\Gamma\lambda_{j}\psi)$ for $\Gamma=i\gamma_{5}$, that have been derived by a similar method [24, 25].
+In this work, the axial and pseudoscalar pion-constituent quark coupling constants will be derived from a quark determinant and investigated by considering nondegenerate up and down quark masses.
+The longstanding problem of establishing relations between quark and hadron level quantities are usually firstly envisaged by means of group theoretical framework although the whole picture might become more intrincated due to several involved issues such as confinement and to the dynamical character of Bethe-Salpeter and Faddeev type equations for bound quarks.
+Although this whole project may depend on the confinment mechanism and related properties, some aspects of hadron are seemingly independent of it according to phenomenology in which the CQM finds support for many hadrons.
+Afterall, color is confined but flavor is not.
+With this program, the validity of the CQM description should be better defined and refined by investigating its consequences such as fine tuning (ISB or FSB) effects in meson couplings.
+Therefore, in the present work, an investigation of the ISB (and FSB) in constituent quark-pion interactions is proposed by means of the pseudoscalar and axial pion couplings to constituent quark currents previously analyzed in the degenerate masses limit [13, 16].
+Although the relation of these quark level couplings to the corresponding nucleon (or baryon in general) is not fully unambiguous for the reasons explicitated above, these quark level calculations and results should lead to more fine and precise assessment of to what extent, why and how constituent quark model works.
 Along these lines, the absolute values of the resulting coupling constants are not exactly the most relevant or interesting point of this work, but rather the effects of the nondegenerate quark masses in each channel, i.e.
@@
 This might contribute to settle a route for analyzing hadron observable world with more complete grounds in low energy QCD.
+As pointed out in [16] momentum dependency of pseudoscalar form factor within this approach is not reasonable due to the momentum independent quark mass.
+Coupling constants, however, are well described.
+Since this level of calculation does not provide the running quark mass, the momentum dependency of form factors will not be investigated in this work.
+The present calculation has a dynamical character since it corresponds to deriving an effective action for pions, being mesons arranged in flavor (and spin) multiplets, and constituent quark currents by articulating the quark determinant obtained from a quark-antiquark interaction mediated by dressed gluons.
+The pion-constituent quark coupling constants will be calculated at different kinematical points from the corresponding dynamical form factors.
 This work is organized as follows.
-In the next section some schematic explanation of the quark determinant in the presence of local meson fields and background quark currents, surrounded by a gluon cloud, is presented.
+In the next section some schematic explanation of the quark determinant, and its large quark mass expansion, in the presence of local meson fields and background quark currents, surrounded by a gluon cloud, is presented.
 The pseudoscalar and axial coupling constants are derived with the corresponding match equations.
@@
 In Section (4) numerical results for the charged and neutral pion pseudoscalar and axial coupling constants are presented for different kinematical points: on shell and off shell (or massless) pions and constituent quarks.
-Some strangeness degree of freedom is shown to provide a small coupling to the neutral pion due to the mixings.
+Some strangeness degree of freedom is shown to provide a small coupling of the strangeness quark currents to the neutral pion due to the flavor mixings.
 In the last Section there is a Summary with conclusions.
 2 Pion couplings from the quark determinant
-The general procedure for the derivation of the quark determinant in the presence of background quark currents, surrounded by a gluon cloud represented by components of a gluon propagator, and local auxiliary meson fields displayed in quark-antiquark flavor U(3) nonets was developed in Refs.
+The general procedure for the derivation of the quark derminant in the presence of background quark currents, surrounded by a gluon cloud represented by components of a gluon propagator, and local auxiliary meson fields displayed in quark-antiquark flavor U(3) nonets was developed in Refs.
 [13, 16, 17].
-By selecting the (chiral) scalar and pseudoscalar auxiliary fields and by performing an usual chiral rotation, the scalar field nonet is eliminated in favor of non-linear parametrization of the (quasi) Goldstone boson octet as fluctuations around the vacuum.
+By selecting the (chiral) scalar and pseudoscalar auxiliary fields and by perfoming an usual chiral rotation, the scalar field nonet is eliminated in favor of non-linear parameterization of the (quasi) Goldstone boson octet as fluctuations around the vacuum.
 It can be written in terms of: $U=e^{\frac{iP\cdot\lambda}{F}}$ and $U^{\dagger}=e^{\frac{-iP\cdot\lambda}{F}}$, where $\lambda_{i}$ are the flavor Gell-Mann matrices and the pseudoscalar meson fields $P_{i}$ have canonical normalization.
@@
 Note that the flavor structure of the positive pion absorption coupling is different from the negative pion absorption but it is the same as the negative pion emission.
-In the expansion of the determinant, there also appear the scalar and vector (two-)pion couplings [13] that are of the same order of magnitude of the pseudoscalar and axial ones, although they present a considerably more intricate isospin structure and they will not be considered in this work.
+In the expansion of the determinant, there also appear the scalar and vector (two-)pion couplings [16] that are of the same order of magnitude of the pseudoscalar and axial ones, although they present a considerably more intrincate isospin structure and they will not be considered in this work.
 Figure 1: Diagrammatic interpretation of the leading (charged and neutral) pion couplings to constituent quark currents (absorption) at one loop level.
@@
 Some complementary definitions that will be useful to work out mixings are presented in the Appendix (A).
-The positive and negative pion form factors present exactly the same form factors since $G_{11}$ is given by Eq.
+Positive and negative pions present exactly the same form factors since $G_{11}$ and $G_{22}$ are given by Eq.
 (A.1).
 In fact, it involves the average of positive or negative pion absorption and emission processes.
+Note that the equations for the axial and pseudoscalar pion couplings are independent of each other in the approach considered in the present work, and none of them is calculated by considering the GTR which, nonetheless, is satisfied quite well when comparing numerical results.
+It is also interesting to emphasize that, as a consequence of the standard methods applied for the derivation of the quark determinant, pion couples directly to the pseudoscalar and axial quark currents.
+It is worth to emphasize that the above form factors were obtained in a dynamical way and not from the QCD vertex function as calculated diagrammatically or in a Schwinger-Dyson equations framework.
+Therefore coupling constants do not emerge as the residue of the pole from the corresponding Bethe-Salpeter equation (BSE) [29, 30].
+The relation between the present approach and the standard one based in BSE has not yet been settled.
 3 Quark currents, meson states and mixings
-By writing interactions (6) explicitly in terms of charge eigenstates, for the pseudoscalar coupling and corresponding pion form factors, it yields:
+By writing interactions (6) explicitely in terms of charge eigenstates, for the pseudoscalar coupling and corresponding pion form factors, it yields:
 $I_{ps}^{B}$ $=$ $G_{11}^{ps}(\pi^{+}\bar{d}i\gamma_{5}u+\pi^{-}\bar{u}i\gamma_{5}d)+G_{33}^{ps}\pi_{3}(\bar{u}i\gamma_{5}u-\bar{d}i\gamma_{5}d),$ (8)
@@
 There are two different possible ways meson mixing can show up.
-First in processes that lead to pions, etas and etaprimes for which contamination from $\eta,\eta^{\prime}$ oscillations can lead to contributions for the neutral pion coupling to constituent quarks/nucleon.
+First in processes that lead to pions, etas and etaprimes for which contaminations from $\eta,\eta^{\prime}$ oscillations can lead to contributions for the neutral pion coupling to constituent quarks/nucleon.
 Second, in processes in which the original neutral pion oscillates to $\eta$ or $\eta^{\prime}$ that can couple to quarks/nucleons.
-Consider the following parametrization of these meson mixings (rotations):
+Consider the following parameterization of these meson mixings (rotations):
 $P_{0}$ $=$ $\bar{G}_{80}\eta+\bar{G}_{03}\pi^{0}+a_{0}\eta^{\prime}$
@@
 where the pion receives the smallest mixings and then $a_{3}<1$ with $a_{3}\sim 1$.
-These equations have been inspired in usual parametrization [26, 27].
+These equations have been inspired in usual parameterizations [31, 32].
 Only two mixing terms will be needed below, $\eta-\pi^{0}$ or $\eta^{\prime}-\pi^{0}$ by means of $\bar{G}_{03}$ and $\bar{G}_{38}$.
@@
 $g_{03}^{\xi}\equiv\bar{G}_{03}^{\xi}=\frac{G_{03}^{\xi}}{2(G_{00}^{\xi}+G_{33}^{\xi})},\;\;\;\;\;g_{38}^{\xi}\equiv\bar{G}_{38}^{\xi}=\frac{G_{38}^{\xi}}{2(G_{33}^{\xi}+G_{88}^{\xi})}.$ (18)
-If these mixings were simply $\bar{G}_{03}^{\xi}=G_{03}^{\xi}$ and $\bar{G}_{38}^{\xi}=G_{38}^{\xi}$ the pion coupling to the strange-quark current would disappear, i.e. $G_{mix,ss}^{\xi}=0$.
+If these mixings were simply $\bar{G}_{03}^{\xi}=G_{03}^{\xi}$ and $\bar{G}_{38}^{\xi}=G_{38}^{\xi}$ the pion coupling to strange quark current would disappear, i.e. $G_{mix,ss}^{\xi}=0$.
 3.2 Renormalization condition
-To provide numerical results, the following renormalization condition will be considered by fixing the pseudoscalar charged pion coupling constant with a value based in phenomenology [28] (off shell pion and on shell quark):
+To provide numerical results, the folllowing renormalization condition will be considered by fixing the pseudoscalar charged pion coupling constant with a value based in phenomenology [33] (off shell pion and on shell quark):
 $G^{ps}_{\pi^{\pm}}(K^{2}=M^{2},Q^{2}=0)=13.$ (19)
 It turns out that, for the sets of parameters considered in this work, results will provide nearly $g_{A}(M^{2},0)^{\pi^{\pm}}\sim 1$ without further conditions.
-A different condition could be imposed by assuming each quark of the nucleon to be responsible for part of the couplings to mesons, such that $G_{ps}\sim 13/3$.
-However, for the condition above, pion-nucleon coupling constants should arise from an average of pion-constituent quark coupling constants.
+Conversely, in other works [16, 34], an alternative condition was considered, i.e.
+to fix either $g_{A}=1$ or $g_{v}=12$ (vector meson coupling to constituent quarks) for the CQM.
+A different condition could be imposed by assuming each quark of the nucleon to be responsible for part of the couplings to mesons along other investigations [35], such that $G_{ps}\sim 13/3$.
+However, for the condition above, pion-nucleon coupling constants should arise from an average of pion-constituent quark coupling constants instead as a sum.
 This might be considered in the Goldberber-Treiman (GT)-relation if the nucleon mass is considered instead of the constituent quark mass ($M_{q}$).
 Otherwise, with a constituent quark mass, GT-relation should be written as $\frac{G^{ps}}{g^{A}}=\frac{M_{q}}{f_{\pi}}(1+\Delta_{GT})$ being $\Delta_{GT}$ not small with respect to 1.
-Conversely, in other works [13, 30], an alternative condition was considered, i.e.
-to fix either $g_{A}=1$ or $g_{v}=12$ (vector meson coupling to constituent quarks) for the CQM.
 4 Numerical Results
@@
 Given that there is an intrinsic ambiguity in defining values for the quark effective masses, results will be shown for different choices of $M_{u},M_{d}$ (and also $M_{s}$).
-These masses, for the chosen gluon effective propagator, determine $G_{ij}^{\xi}$ (both $G_{ps}$ and $G_{A}$) numerically, and the different sets of parameters (s.o.p.) are shown in Table (1).
-Sets of parameters $A_{1}$ and $A_{2}$ were extracted from [21].
-The third $A_{3}$ is based in often values found in the literature.
+These masses, for the chosen gluon effective propagator, determine $G_{ij}^{\xi}$ (both $G^{ps}$ and $G^{A}$) numerically, and the different sets of parameters (s.o.p.) are shown in Table (1).
+In the present work, we consider quark effective masses - sets of parameters $A_{1}$ and $A_{2}$ obtained from a (flavor-dependent) NJL model as presented in [24].
+In NJL-type models there is need to fix chosen observables/quantities to be mass independent and this choice may bring differences in the results, for example in Refs.
+[38], being the neutral and charged pion masses chosen to fit the model parameters.
+In the present calculation however, the renormalization condition presented in Eq.
+(19) is a further "constraint" for the resulting effective action and numerical results.
+The third s.o.p., $A_{3}$, is based in values often found in the literature.
 Table 1: Sets of parameters for quark effective masses.
@@
 In Table (3) values of pseudoscalar coupling constants $G_{ps}$, defined in different kinematical points (off shell or on shell constituent quark and pion) are shown for the different sets of parameters presented in Table (1).
-Besides the charged pion coupling constant, the neutral pion coupling constant is exhibited with and without flavor states mixings, respectively $G^{ps}_{\pi^{0}}(0)$ and $G^{ps}_{\pi^{0}}(ij)$ from Eqs.
+The following kinematical points have been considered:
+$S_{1}:(K^{2},Q^{2})=(M_{f}^{2},0),$ $\;\;\;S_{2}:(K^{2},Q^{2})=(M_{f}^{2},M_{\pi}^{2}),$ (22)
+$S_{3}:(K^{2},Q^{2})=(0,0),$ $\;\;\;S_{4}:(K^{2},Q^{2})=(0,M_{\pi}^{2}).$
+Besides the charged pion coupling constant, the neutral pion coupling constant is exhibitted with and without flavor states mixings, respectively $G^{ps}_{\pi^{0}}(0)$ and $G^{ps}_{\pi^{0}}(ij)$ from Eqs.
 (11) and (17).
 As shown above the neutral pion coupling can be decomposed in the couplings of states $\bar{u}u$ and $\bar{d}d$ states, that might also receive contributions from flavor state mixings - and these are shown in the middle part of this Table.
-In the last two columns the pseudoscalar coupling constant of the neutral pion coupling to strange quark current $G_{mix,ss}^{ps}$ is exhibited with a parameter $x_{ps}^{m}$ that will be defined below.
+In the last two columns the pseudoscalar coupling constant of the neutral pion coupling to strange quark current $G_{mix,ss}^{ps}$ is exhibitted with a parameter $x_{ps}^{m}$ that will be defined below.
 Results for s.o.p.
-$A_{2}$ ($A_{1}$), with the lower (larger) values of quark masses, present the smallest (largest) variation of values for the different kinematical points.
+$A_{2}$ ($A_{1}$), with the lower (larger) values of quark masses, present the smallest (largest) variation of values for the different kinetical points.
 The charged pion coupling constants are smaller than the neutral pion ones except for the case of the fourth kinematical point: off shell quarks and on shell pions $(K^{2}=0,Q^{2}=M_{\pi}^{2})$.
 The neutral pion coupling to the $\bar{u}u$ quark current has larger strength than the coupling to $\bar{d}d$ for all kinematical points, and for both cases, without and with mixings, $G^{ps}_{uu}(0)$ and $G^{ps}_{uu}(ij)$ respectively.
-Recent analysis of the pion-nucleon pseudoscalar coupling constant calculated with different approaches [6, 5] exhibit a very similar trend with small numerical difference between charged and neutral pion couplings to nucleons.
-The neutral pion -nucleon coupling constants are slightly larger than the charged pion coupling coupling constant.
-For the sake of comparison of the behavior, the following values can be extracted from the calculation with the EFT approach [6]: $g_{\pi^{0}pp}=13.22688$, $g_{\pi^{0}nn}=13.39731$, $g_{\pi^{\pm}np}=13.22611$.
-According to the quark model, proton and neutron structures suggest that the corresponding coupling to the pion might be a sort of sum or average for the pion coupling to corresponding states, $p\sim uud$ and $n\sim ddu$ respectively.
-However, due to the choice of the renormalization condition ($G^{ps}_{\pi^{p}m}=13$), the resulting neutral pion coupling to nucleons, as a first simplified approach, might be considered as a simple average:
+Recent analysis were done of the pion-nucleon pseudoscalar coupling constant with different approaches [6, 7].
+With different definitions of coupling constants, usually denoted by $f_{\pi^{a}NN^{\prime}}$ for $a=0,\pm$ and $N,N^{\prime}=n,p$, results exhibit a somehow similar trend with small numerical difference between charged and neutral pion couplings to nucleons.
+For the sake of comparison of the behavior, some values are reproduced in Table (2).
+The first one is the value extracted from Ref.
+[7], the second and third lines correspond to values presenting smallest standard deviation confidence level from Ref.
+[6].
+In spite of the differences in absolute values the analysis will be focused on the relative values.
+Whereas the coupling constant of neutral pion to neutrons is larger than the coupling constant to protons in both approaches, there is a considerable difference between the results of these approaches concerning the charged pion coupling to the nucleon.
+Table 2: Neutral and charged pion-nucleon pseudoscalar coupling constants extracted from different recent works [6, 7].
+$G_{\pi^{0}nn}$ $G_{\pi^{0}pp}$ $G_{\pi^{\pm}np}$
+$1$- [7] 13.397 13.227 13.226
+$2$- [6] 15.455 12.511 18.571
+$3$- [6] 13.452 13.174 18.571
+According to most versions of the quark model, proton and neutron couple to mesons by means of quarks in an addictive way [15, 35].
+Along this idea, the pion coupling to nucleons might also arise in a sort of average for the pion coupling to corresponding quark currents, $p\sim uud$ and $n\sim ddu$.
+Due to the choice of the renormalization condition ($G^{ps}_{\pi^{\pm}}=13$), the resulting neutral pion coupling to nucleons, as a first simplified approach, might be considered as a simple average:
 $g_{\pi^{0}nn}\sim\frac{2G_{\pi^{0}dd}+G_{\pi^{0}uu}}{3},\;\;\;\;\;\;\;\;g_{\pi^{0}pp}\sim\frac{2G_{\pi^{0}uu}+G_{\pi^{0}dd}}{3}.$ (23)
-Concerning the charged pion coupling, the behavior shown in the Table is consistent with the results for the charged pion-neutron-proton and neutral pion-proton coupling constants such that that is $g_{\pi^{\pm}np}<g_{\pi^{0}pp},g_{\pi^{0}nn}$ and $g_{\pi^{\pm}np}\sim g_{\pi^{0}pp}$ [6, 5].
-Along these lines, it is also expected that $g_{\pi^{0}nn}>g_{\pi^{0}pp}$.
-This hierarchy in pion-nucleon coupling constants, however, does not happen with the values shown in the Table for (22) for the model of (22), since it would require $G_{\pi^{0}dd}>G_{\pi^{0}uu}$ i.e.
-the opposite behavior to the one shown in the Table.
-It is shown below that mixings can mend this behavior.
-By varying the strength of the mixing interactions by a constant uniform factor $x^{m}_{ps}$, the pseudoscalar coupling constants $G^{ps}_{uu}$ and $G^{ps}_{dd}$ can flip their hierarchy if $x_{ps}^{m}<0$.
+The behavior shown in the Table (3) is partially consistent with the results for the charged pion-neutron-proton and neutral pion-neutron coupling constants of Ref.
+[6] by considering these averages, such that that is $g_{\pi^{\pm}np}>g_{\pi^{0}nn}$.
+However, the behavior of values in Table (3) is rather similar to [7] when comparing charged pion to neutral-pion proton coupling constants, i.e.
+$g_{\pi^{\pm}np}\sim g_{\pi^{0}pp}$ and $g_{\pi^{\pm}np}<g_{\pi^{0}pp}$.
+For the results shown in the Table (3) it would be expected that $g_{\pi^{0}nn}<g_{\pi^{0}pp}$ and this is not compatible with the results shown in Table (2).
+Overall, there is only partially agreement between results of these two quoted works and the present one.
+It is shown below that mixings can modify the relative behavior of $G_{uu}^{ps}(ij)$ and $G_{dd}^{ps}(ij)$, and then $g_{\pi^{0}nn}$ and $g_{\pi^{0}pp}$.
+NJL-type models, have been considered in [39] where an hierarchy of coupling constant values has been obtained that is (at least partially) in disagreement with all these quoted works and with the present one.
+By varying the strength of the mixing interactions by an arbitrary constant uniform factor $x^{m}_{ps}$, the pseudoscalar coupling constants $G^{ps}_{uu}$ and $G^{ps}_{dd}$ can flip their hierarchy (suitable for making $g_{\pi^{0}pp}<g_{\pi^{0}nn}$) if $x_{ps}^{m}<0$.
 Therefore mixing parameters were redefined by:
@@
 (24).
-s.o.p./$(K^{2},Q^{2})$ $G^{ps}_{\pi^{\pm}}$ $G^{ps}_{\pi^{0}}$ (0) $G^{ps}_{\pi^{0}}$ ($ij$) $G^{ps}_{uu}$ (0) $G^{ps}_{uu}$ ($ij$) $G^{ps}_{dd}$ (0) $G^{ps}_{dd}$ ($ij$) $G^{ps}_{mix,ss}$ $x^{m}_{ps}$
+s.o.p. $G^{ps}_{\pi^{\pm}}$ $G^{ps}_{\pi^{0}}$ (0) $G^{ps}_{\pi^{0}}$ ($ij$) $G^{ps}_{uu}$ (0) $G^{ps}_{uu}$ ($ij$) $G^{ps}_{dd}$ (0) $G^{ps}_{dd}$ ($ij$) $G^{ps,ss}_{mix}$ $x^{m}_{ps}$
 $A_{1}$
-$(M_{f}^{2},0)$ 13.000 13.001 13.007 13.471 13.648 12.531 12.367 0.05 -1.7
-$(M_{f}^{2},M_{\pi}^{2})$ 10.446 10.546 10.553 11.098 11.244 9.994 9.863 0.37 -1.2
-$(0,0)$ 17.870 17.871 17.875 18.172 18.411 17.569 17.338 - 0.01 -2.3
-$(0,M_{\pi}^{2})$ 18.106 18.091 18.095 18.399 18.642 17.782 17.545 -0.01 -2.3
+$S_{1}$ 13.000(1) 13.001(2) 13.007(2) 13.471(1) 13.648(1) 12.531(1) 12.367(1) 0.05(1) -1.7(2)
+$S_{2}$ 10.446(1) 10.546(2) 10.553(2) 11.098(1) 11.244(1) 9.994(1) 9.863(1) 0.37(1) -1.2(2)
+$S_{3}$ 17.870(1) 17.871(2) 17.875(2) 18.172(1) 18.411(1) 17.569(1) 17.338(1) - 0.01(1) -2.3(2)
+$S_{4}$ 18.106(1) 18.091(2) 18.095(2) 18.399(1) 18.642(1) 17.782(1) 17.545(1) -0.01(1) -2.3(2)
 $A_{2}$
-$(M_{f}^{2},0)$ 13.000 13.000 13.003 13.363 13.463 12.637 12.543 -0.02 -2.6
-$(M_{f}^{2},M_{\pi}^{2})$ 12.263 12.284 12.287 12.708 12.802 11.861 11.773 -0.02 -2.5
-$(0,0)$ 14.550 15.550 14.552 14.818 14.928 14.283 14.177 -0.01 -2.6
-$(0,M_{\pi}^{2})$ 14.808 14.792 14.794 15.069 15.181 14.514 14.406 -0.01 -2.2
+$S_{1}$ 13.000(1) 13.000(2) 13.003(2) 13.363(1) 13.463(1) 12.637(1) 12.543(1) -0.02(1) -2.6(2)
+$S_{2}$ 12.263(1) 12.284(2) 12.287(2) 12.708(1) 12.802(1) 11.861(1) 11.773(1) -0.02(1) -2.5(2)
+$S_{3}$ 14.550(1) 15.550(2) 14.552(2) 14.818(1) 14.928(1) 14.283(1) 14.177(1) -0.01(1) -2.6(2)
+$S_{4}$ 14.808(1) 14.792(2) 14.794(2) 15.069(1) 15.181(1) 14.514(1) 14.406(1) -0.01(1) -2.2(2)
 $A_{3}$
-$\;(M_{f}^{2},0)$ 13.000 13.000 13.004 13.352 13.473 12.649 12.535 0.004 -2.0
-$(M_{f}^{2},M_{\pi}^{2})$ 11.735 11.779 11.783 12.191 12.301 11.368 11.266 0.04 -1.8
-$(0,0)$ 15.458 15.459 15.461 15.704 15.846 15.213 15.076 -0.01 -2.3
-$(0,M_{\pi}^{2})$ 15.702 15.686 15.689 15.940 16.083 15.433 15.294 -0.01 -2.3
+$S_{1}$ 13.000(1) 13.000(2) 13.005(2) 13.352(1) 13.537(1) 12.649(1) 12.474(1) - 0.004(5) -2.0(2)
+$S_{2}$ 11.735(1) 11.779(2) 11.783(2) 12.191(1) 12.301(1) 11.368(1) 11.266(1) 0.04(1) -1.8(2)
+$S_{3}$ 15.458(1) 15.459(2) 15.461(2) 15.704(1) 15.846(1) 15.213(1) 15.076(1) -0.01(1) -2.3(2)
+$S_{4}$ 15.702(1) 15.686(2) 15.689(2) 15.940(1) 16.083(1) 15.433(1) 15.294(1) -0.01(1) -2.3(2)
 Table 4: Axial pion coupling constants for charged and neutral pions at different kinematical points, without ($G^{A}_{\pi^{0}}(0)$) and with ($G^{A}_{\pi^{0}}(ij)$) mixing interactions $G_{i\neq j}$, similarly to Eqs.
@@
 In the last column the axial coupling constant of neutral pion -strange quark current.
-s.o.p./$(K^{2},Q^{2})$ $G^{A}_{\pi^{\pm}}$ $G^{A}_{\pi^{0}}$ (0) $G^{A}_{\pi^{0}}$ ($ij$) $G^{A}_{uu}$ (0) $G^{A}_{uu}$ ($ij$) $G^{A}_{dd}$ (0) $G^{A}_{dd}$ ($ij$) $G^{A}_{mix,ss}$
+s.o.p. $G^{A}_{\pi^{\pm}}$ $G^{A}_{\pi^{0}}$ (0) $G^{A}_{\pi^{0}}$ ($ij$) $G^{A}_{uu}$ (0) $G^{A}_{uu}$ ($ij$) $G^{A}_{dd}$ (0) $G^{A}_{dd}$ ($ij$) $G^{A}_{mix,ss}$
 $A_{1}$
-$(M_{f}^{2},0)$ 1.6027 1.6026 1.6026 1.6000 1.5954 1.6051 1.6099 0.0001
-$(M_{f}^{2},M_{\pi}^{2})$ 1.8888 1.8750 1.8750 1.8717 1.8662 1.8784 1.8839 0.0002
-$(0,0)$ 1.1936 1.1936 1.1935 1.1970 1.1935 1.1901 1.1936 - 0.0002
-$(0,M_{\pi}^{2})$ 1.2230 1.2210 1.2210 1.2250 1.2214 1.2170 1.2206 - 0.0002
+$S_{1}$ 1.603(1) 1.603(2) 1.603(2) 1.600(1) 1.595(1) 1.605(1) 1.610(1) 0.001(1)
+$S_{2}$ 1.889(1) 1.875(2) 1.875(2) 1.872(1) 1.866(1) 1.878(1) 1.8834(1) 0.001(1)
+$S_{3}$ 1.194(1) 1.194(2) 1.194(2) 1.197(1) 1.194(1) 1.190(1) 1.194(1) - 0.001(1)
+$S_{4}$ 1.223(1) 1.221(2) 1.221(2) 1.225(1) 1.221(1) 1.217(1) 1.221(1) - 0.001(1)
 $A_{2}$
-$(M_{f}^{2},0)$ 0.9736 0.9735 0.9736 0.9706 0.9656 0.9765 0.9815 0.0002
-$(M_{f}^{2},M_{\pi}^{2})$ 1.1442 1.1357 1.1357 1.1320 1.1262 1.1394 1.1452 0.0002
-$(0,0)$ 0.7801 0.7801 0.7801 0.7819 0.7779 0.7782 0.7822 -0.0001
-$(0,M_{\pi}^{2})$ 0.8062 0.8045 0.8045 0.8070 0.8029 0.8019 0.8060 -0.0001
+$S_{1}$ 0.974(1) 0.974(2) 0.974(2) 0.971(1) 0.966(1) 0.976(1) 0.982(1) 0.001(1)
+$S_{2}$ 1.144(1) 1.136(2) 1.136(2) 1.132(1) 1.126(1) 1.139(1) 1.145(1) 0.001(1)
+$S_{3}$ 0.7801(1) 0.780(2) 0.780(2) 0.782(1) 0.778(1) 0.778(1) 0.782(1) -0.001(1)
+$S_{4}$ 0.806(1) 0.804(2) 0.804(2) 0.807(1) 0.803(1) 0.802(1) 0.806(1) -0.001(1)
 $A_{3}$
-$(M_{f}^{2},0)$ 1.1602 1.1601 1.1601 1.1577 1.1536 1.1626 1.1666 0.0001
-$(M_{f}^{2},M_{\pi}^{2})$ 1.3647 1.3547 1.3547 1.3517 1.3469 1.3578 1.3626 0.0002
-$(0,0)$ 0.9030 0.9030 0.9029 0.9052 0.9020 0.9008 0.9040 -0.0001
-$(0,M_{\pi}^{2})$ 0.9296 0.9278 0.9278 0.9305 0.9273 0.9251 0.9283 -0.0001
-As noted before, the positive pion absorption (negative pion emission) coupling constant in both channels, $G_{du}^{\xi}(K^{2},Q^{2})$ ($\xi=ps,A$), might be different from the negative pion absorption (positive pion emission) coupling constant, $G_{ud}^{\xi}(K^{2},Q^{2})$ in the case of non-degenerate quark masses.
+$S_{1}$ 1.160(1) 1.160(2) 1.160(2) 1.158(1) 1.154(1) 1.163(1) 1.167(1) 0.001(1)
+$S_{2}$ 1.365(1) 1.355(2) 1.355(2) 1.352(1) 1.347(1) 1.358(1) 1.363(1) 0.001(1)
+$S_{3}$ 0.903(1) 0.903(2) 0.903(2) 0.905(1) 0.902(1) 0.901(1) 0.904(1) -0.001(1)
+$S_{4}$ 0.930(1) 0.928(2) 0.928(2) 0.931(1) 0.927(1) 0.925(1) 0.928(1) -0.001(1)
+As noted before, the positive pion absorption (negative pion emission) coupling constant in both channels, $G_{du}^{\xi}(K^{2},Q^{2})$ ($\xi=ps,A$), might be different from the negative pion absorption (positive pion emission) coupling constant, $G_{ud}^{\xi}(K^{2},Q^{2})$ in the case of nondegenerate quark masses.
 The difference between these two charged pion coupling constants, for both pseudoscalar and axial couplings, is shown in Fig.
@@
 $D_{u-d}\equiv\frac{G_{ud}^{\xi}(K^{2}=M_{f}^{2},Q^{2})-G_{du}^{\xi}(K^{2}=M_{f}^{2},Q^{2})}{G_{ud}^{\xi}(K^{2}=M_{f}^{2},Q^{2})+G_{du}^{\xi}(K^{2}=M_{f}^{2},Q^{2})}.$ (25)
-Two cases are exhibited, $Q^{2}=M_{\pi^{\pm}}^{2}$ and $Q^{2}=0$.
+Two cases are exhibitted, $Q^{2}=M_{\pi^{\pm}}^{2}$ and $Q^{2}=0$.
 It is seen that the variation is very small and basically linear, i.e.
@@
 5 Summary and conclusions
-Pseudoscalar and axial charged and neutral pion -constituent quark coupling constants were derived and investigated by considering quark mass non-degeneracy within an approach proposed previously by the author.
-These coupling constants were calculated for different sets of parameters of constituent quark masses and in different kinematical points: for on shell or off shell pions and constituent quarks.
+Pseudoscalar and axial charged and neutral pion -constituent quark coupling constants were derived and investigated by considering quark mass non-degeneracy within a dynamical approach proposed previously by the author.
+In this approach, coupling constants and form factors are obtained from a quark determinant in the presence of meson fields and quark currents.
+Gluon effects are considered by means of a gluon effective propagator that dresses background quark currents leading to constituent quark currents.
+These coupling constants were calculated numerically for different sets of parameters of constituent quark masses and in different kinematical points: for on shell or off shell pions and constituent quarks.
 The pseudoscalar coupling constant calculated for off shell (or massless) charged pion and on shell constituent quark was fixed as a renormalization condition by identifying it to a phenomenological value $G^{ps}_{\pi^{\pm}}(K^{2}=M_{f}^{2},Q^{2}=0)=13$.
@@
 $g_{A}\sim 1$.
-Being all the coupling constants functions of the quark constituent masses (and not the small current quark masses), the Goldberger-Treiman relation at the quark level presents, in fact, a large correction for the values $G_{ps}=13$ and $g_{A}=1$.
+Being all the coupling constants functions of the quark constituent masses (and not the small current quark masses), the Goldberger Treiman relation - as usually defined - at the quark level may present, in fact, a large correction for the values $G_{ps}=13$ and $g_{A}=1$.
+The differences between charged and neutral pion couplings to the nucleon recently obtained in different approaches are small and this trend is reproduced with the constituent quark couplings in this work.
+The specific hierarchies of the different coupling constants, however, are different among each other, from [6, 7, 39] and also the present.
 Results suggest that lower values of quark effective masses lead to smaller differences between the neutral and charged pion couplings (both axial and pseudoscalar).
-A very small difference between charged and neutral pion couplings has been noted in the pion-nucleon pseudoscalar coupling constant.
-The neutral pion coupling to quark currents $\bar{u}u$ and $\bar{d}d$ was compared to neutral pion coupling to protons and neutrons (respectively) and this comparison suggests mixing interactions stronger than those provided by flavor symmetry breaking are needed.
-This could be provided by instanton induced ’t Hooft-like interactions.
-The resulting charged pion coupling constants from the quark determinant present full U(1) and CP invariance with $G_{11}=G_{22}$.
+The average of neutral pion couplings to quark currents $\bar{u}u$ and $\bar{d}d$ were compared to neutral pion coupling to protons and neutrons and this comparison suggests mixing interactions should be stronger than those provided by flavor symmetry breaking to reproduce values from other works such as Refs.[6, 7].
+This extra mixing could be provided by instanton induced ’t Hooft-like interactions.
+A more extensive investigation of the role of the specific gluon propagator will be provided elsehwere.
+The resulting charged pion coupling constants from the quark determinant present full U(1) and CP invariances with $G_{11}=G_{22}$.
 It corresponds to an average of emitted and absorbed (positive or negative) pion, being therefore the same for each charge eigenstate or for each type of process (emission or absorption).
@@
 However, high precision experimental results/fittings would be needed to identify such difference.
-Interactions that mix neutral flavor eigenstates $i,j=0,3,8$, leading to usual $\pi^{0}-\eta-\eta^{\prime}$ mixing, have been taken into account for both the quark currents and for the neutral pion.
+FSB induced interactions, that mix neutral flavor eigenstates, $i,j=0,3,8$, leading to usual $\pi^{0}-\eta-\eta^{\prime}$ mixing, have been taken into account for both the quark currents and for the neutral pion.
 The shifts in the neutral pion coupling to the quark current $\bar{u}u$ are opposite to the shifts in the coupling to the current $\bar{d}d$.
 Besides that, a quite weak neutral pion coupling to the strange current $\bar{s}s$ was obtained for both axial and pseudoscalar cases.
-The relative strength of these couplings to strangeness, with respect to the pseudoscalar and axial ones, is very small $\sim 10^{-3},10^{-4}$, so that their contributions for neutral pion couplings to strange baryons would receive very tiny shifts only possibly identified with high precision measurements.
-Although the hierarchy of the strength of the flavor dependent coupling constants might nearly follow the one from the nucleon-pion couplings, the behavior of these coupling constants by inserting these couplings into the three-quark bound system for the nucleon can bring further insights into this subject.
-Besides that, at finite temperature, baryon density or external electromagnetic fields, all these couplings must receive non-negligible contributions.
+The relative strength of these couplings to strange current with respect to the pseudoscalar and axial ones, is very small $\sim 10^{-3},10^{-4}$, so that their contributions for neutral pion couplings to strange baryons would receive very tiny shifts only possibly identified with high precision measurements.
+However, note that this coupling is proportional to the mixing interaction $G_{i\neq j}$ that must receive contributions from ’t Hooft-type interaction and sixth order interactions and these might increase considerably its value.
+Nevertheless, given that hadrons are strongly interacting systems, small contributions might, possibly, lead to sizeable effects.
+Although the strength of the flavor dependent constituent quark coupling constants might follow an hierarchy the behavior of these coupling constants by inserting these couplings into the three-quark bound system for the nucleon can bring further insights into this subject.
+Besides that, at finite temperature, baryon density or external electromagnetic fields, all these couplings must receive non-negligeable contributions.
 This type of investigation might seem not appropriated for the understanding of hadron interactions because quarks are confined in hadrons.
-However it is important to emphasize that confinement operates in color degrees of freedom whereas flavor (isospin) cannot be (completely) confined, otherwise all the quark-based description of hadrons, and maybe QCD, would be invalidated.
+However it is important to emphasize that confinement operates in color degrees of freedom whereas flavor-isospin cannot be (completely) confined, otherwise all the quark-based description of hadrons, and maybe QCD, would be invalidated.
 Acknowledgements
-The author thanks a F.L.B. is a member of INCT-FNA, Proc. 464898/2014-5 and he acknowledges partial support from CNPq-312750/2021 and CNPq-407162/2023-2.
+The author thanks short discussion with C.D.
+Roberts in early stages of this work.
+F.L.B.
+is a member of INCT-FNA, Proc.
+464898/2014-5 and he acknowledges partial support from CNPq-312750/2021-8 and CNPq-407162/2023-2.
 Appendix A Some relations